# Hopf formulae for homology of skew braces

### v1

@@ 1. Introduction
Skew braces form a variety of universal algebras $\mathsf{SKB}$, and generalise at the same time groups and radical rings. Concretely, given a group $(G,\cdot)$ one can give $G$ a skew brace structure by taking +=+$\cdot$+ = ⋅ and $\circ$=$\cdot$ = ⋅. In particular, this means that the variety $\mathsf{Grp}$ of groups is a subvariety of the variety $\mathsf{SKB}$ of skew braces. On the other hand, a radical ring $(R,+,\cdot)$ is a (not necessarily commutative) ring without unit such that the operation $a\circ b=a+a\cdot b+b$ is a group operation. In this case, the triple $(R,+,\circ)$ is a skew brace. We will recall in Section 4 a characterisation of radical rings [25] that implies that radical rings also form a subvariety $\mathsf{RadRng}$ of the variety $\mathsf{SKB}$ of skew braces. Since skew braces form a variety of $\Omega$-groups in the sense of Higgins [14], hence in particular a semi-abelian category [19] (in fact, even a strongly protomodular category [2, 3]), it is natural to look at non-abelian homology of skew braces with coefficient functors in the subvarieties $\mathsf{Grp}$ of groups, $\mathsf{Ab}$ of abelian groups, and $\mathsf{RadRng}$ of radical rings.
@@
via the classical Hopf formula, i.e. the right-hand isomorphism. The commutators $[K,F]$ and $[F,F]$ appearing in this formula are “relative” to the chosen subvariety of the variety $\mathsf{Grp}$ of groups, in this case the variety $\mathsf{Ab}$ of abelian groups, so that they could also be denoted by $[K,F]_{\mathsf{Ab}}$ and $[F,F]_{\mathsf{Ab}}$, respectively. The homology group $\mathsf{H}_{2}(G)$ can also be obtained by applying comonadic homology theory [1] to the variety $\mathsf{Grp}$ with respect to the abelianisation functor $\mathsf{ab}\colon\mathsf{Grp}\rightarrow\mathsf{Ab}$, so that $\mathsf{H}_{2}(G)\cong\mathsf{H}_{2}(G,\mathsf{ab})$, the right-hand side homology group now being the one arising from the comonadic approach by taking $\mathsf{ab}$ as coefficient functor:
@@
First, in Section 3, we characterize the central extensions in $\mathsf{SKB}$ corresponding to the subvariety $\mathsf{Grp}$ of groups (Proposition 3.5) that are the surjective homomorphisms $f\colon A\rightarrow B$ of skew braces such that $[\operatorname{Ker}(f),A]_{\mathsf{Grp}}=0$, where $[\operatorname{Ker}(f),A]_{\mathsf{Grp}}$ is the additive subgroup of $A$ generated by all the elements of the form
@@
Section 4 is devoted to the study of central extensions of skew braces relative to the subvariety $\mathsf{RadRng}$ of radical rings. For this purpose we introduce a new object, the radicalator (Definition 4.9). This is generated as a subgroup of the additive group by the elements $(a+b)\circ c-b\circ c+c-a\circ c$ and $a+b-a-b$ for all $a,b,c\in A$. The radicalator turns out to be an ideal $[A,A]_{\mathsf{RadRng}}$ (Proposition 4.8) playing a similar role with respect to the reflection from $\mathsf{SKB}$ to $\mathsf{RadRng}$ to the role played by the derived subgroup in the reflection from $\mathsf{Grp}$ to $\mathsf{Ab}$. Indeed, the quotient $A\rightarrow\frac{A}{[A,A]_{\mathsf{RadRng}}}$ is the universal one transforming a skew brace $A$ into a radical ring $\frac{A}{[A,A]_{\mathsf{RadRng}}}$, and an extension $f\colon A\rightarrow B$ of skew braces is central relatively to the subvariety $\mathsf{RadRng}$ if and only if $[\operatorname{Ker}(f),A]_{\mathsf{RadRng}}=0$ (Theorem 4.13). This pushes forward Rump’s connection between skew braces and radical rings [25] (see also [20]). We also observe that the central extensions of skew braces relative to the subvariety of braces (= skew braces of abelian type) can also be characterized in algebraic terms (Remark 4.14).
@@
is shown to be an ideal (Proposition 5.3), denoted by $[I,A]_{\mathsf{Ab}}$ (here $[a,b]_{+}$ and $[a,b]_{\circ}$ are the usual group-theoretic commutators with respect to the operations $+$ and $\circ$, respectively). An extension $f\colon A\rightarrow B$ of skew braces is then central with respect to $\mathsf{Ab}$ if and only if $[\operatorname{Ker}(f),A]_{\mathsf{Ab}}=0$ (Proposition 5.4). Note that these central extensions of skew braces have already been studied in [23, 24, 21]. We also observe that the commutator $[\operatorname{Ker}(f),A]_{\mathsf{Ab}}$ coincides with the Huq commutator defined in [15] (Proposition 5.5).
@@ 2. Preliminaries
We can introduce a binary operation $*\colon A\times A\rightarrow A$ defined by $a*b=-a+a\circ b-b$ for all $a,b\in A$. By definition, a normal subgroup $I$ of $A_{+}$ is an ideal of $A$ if $a*x\in I$ and $x*a\in I$ for all $a\in A$ and $x\in I$. A normal subgroup $I$ of of the group $(A,+)$ such that $I$ is a subgroup of $A_{\circ}$ and $A*I\subset I$ is called a strong left ideal.
@@
(2.2) ${{A\times_{B}A}}$${A}$${A}$${B}$$\scriptstyle{t}$$\scriptstyle{s}$$\scriptstyle{f}$$\scriptstyle{f}$
where $A\times_{B}A=\{(a_{1},a_{2})\in A\times A\,\mid\,f(a)=f(b)\}$, while $s$ and $t$ are the first and second projection, respectively. The definition of central extension relative to a given subvariety $\mathcal{X}$ of $\mathsf{SKB}$ (as in the work of the Fröhlich school [11, 12, 22]) is then the following:
@@ 3. The subvariety of groups
where $\mathsf{Grp}$ is the subvariety of $\mathsf{SKB}$ whose objects are skew braces with the property that the two group operations $+$ and $\circ$ are equal. This subvariety is isomorphic to the category of groups, of course, and this justifies the slight abuse of notation. The functor $U_{G}$ is a full inclusion, and in order to describe its left adjoint $F$ in (3.1) the following definition will be needed:
@@
Let $A$ be a skew brace, we denote by $A*A$ the subgroup of $A_{+}$ generated by the elements $a*b$ for all $a,b\in A$.
It is well known that $A*A$ is an ideal of $A$ and it is the smallest ideal $I$ such that the quotient $A/I$ is a group.
In our situation the reflector $F:\mathsf{SKB}\to\mathsf{Grp}$ is the functor sending the skew brace $A$ to the quotient $F(A)=A/(A*A)$ by the ideal $A*A$ (with obvious definition on morphisms). Note that here the ideal $A*A$ is precisely the kernel $R(A)$ of the $A$-component of the unit $\eta_{A}\colon A\rightarrow A/(A*A)$ of the adjunction (3.1). This adjunction is admissible in the sense of Categorical Galois theory, since $\mathsf{Grp}$ is a subvariety of the variety $\mathsf{SKB}$ (see Section $5$ in [18]).
@@
Let $A$ be a skew brace and $I\subset A$ an ideal, we define $[I,A]_{\mathsf{Grp}}$ to be the additive subgroup of $A$ generated by the set
@@
Let $A$ be a skew brace and $I\subset A$ an ideal of $A$. Then $[I,A]_{\mathsf{Grp}}$ is an ideal.
@@
An extension $(A,f)$ is central with respect to the adjunction (3.1) if and only if $[\operatorname{Ker}(f),A]_{\mathsf{Grp}}=0$.
@@
Since $(A,f)$ is central, by definition $a*b=R(s)(a*b,0)=R(t)(a*b,0)=0$. Similarly, one can show that $b*a=0$. As $b\in A$ and $a\in\operatorname{Ker}(f)$ were chosen arbitrarily, it follows that $[\operatorname{Ker}(f),A]_{\mathsf{Grp}}=0$.
Assume now that $[\operatorname{Ker}(f),A]_{\mathsf{Grp}}=0$. Let $(a,b)$ and $(a_{1},b_{1})$ be two elements in $A\times_{B}A$. Since $-a_{1}+b_{1}\in\operatorname{Ker}(f)$ one has that $a_{1}\circ(-a_{1}+b_{1})=b_{1}$ so that $-a_{1}+b_{1}=a_{1}^{\prime}\circ b_{1}$. It follows that
@@ 4. The subvariety of radical rings
$[a,b,c]=(a+b)\circ c-b\circ c+c-a\circ c.$
@@
Let $A$ be a skew brace, and $I\subset A$ an ideal. We define $[I,A]_{\mathsf{RadRng}}$ to be the additive subgroup of $A$ generated by the elements $[a,b,c]$, $[c,a,b]$, $[b,c,a]$ and $a+b-a-b$ for all $a\in I$ and $b,c\in A$.
Our next goal is to show that $[I,A]_{\mathsf{RadRng}}$ is an ideal.
@@
If we write $J=[I,A]_{\mathsf{RadRng}}$, then $J_{+}$ is a normal subgroup of $A_{+}$. Indeed, for any $j\in J\subset I$ and $a\in A$, the element $j+a-j-a$ belongs to $J$, so that $J$ is stable under conjugation in $A$.
@@
To see that $\lambda_{a}(J)\subset J$, it suffices to check on the generators of $J$. It is clear that $\lambda_{d}([I_{+},A_{+}])\subset[I_{+},A_{+}]\subset J$ for all $d\in A$. Let then $a,b,c,d\in A$, we must show that $\overline{\lambda_{d}([a,b,c])}=0$ in the quotient group $A_{+}/J_{+}$ where $a,b,c\in A$ and at least one of these elements is in $I$.
@@
$=\overline{d\circ a\circ c}-\overline{c}+\overline{\lambda_{d}(b)\circ c}-\overline{d\circ b\circ c}+\overline{d\circ c}-\overline{d\circ a\circ c}$
$=\overline{d\circ a\circ c}-\overline{c}+\overline{c-d\circ c+d\circ b\circ c}-\overline{d\circ b\circ c}+\overline{d\circ c}-\overline{d\circ a\circ c}$
@@
$-c+\lambda_{d}(b)\circ c\in I,\quad-d\circ b\circ c+d\circ c=\lambda_{d}(-b\circ c+c)\in I.$
@@
Let $A$ be a skew brace, and $I\subset A$ an ideal. The subgroup $[I,A]_{\mathsf{RadRng}}$ is an ideal of $A$.
@@
Let $J=[I,A]_{\mathsf{RadRng}}$. It is left to show that $J*A\subset J$. It suffices to show that $a\circ b-b\in J$ for all $a\in J$ and $b\in A$. Since $J\subset I$, it is sufficient to show this on the generators of $J$. Let $a\in J$ and $b,c\in A$. By Lemma 4.5,
@@
To obtain the second and third equality, we used Corollary 4.6. Assume now that $b\in J$ and $a,c,d\in A$. (This situation is similar to the previous one, except that now we use the fact that multiplying and subtracting on the right by $d$ preserves elements of the commutator subgroup $[I_{+},A_{+}]$.) Then
@@
Finally, assume that $c\in J$ and $a,b,d\in A$. In this case,
@@
Therefore $[k^{\prime}\circ k_{1},b,c]=0$. Similarly one sees that $[b,k^{\prime}\circ k_{1},c]=0$. Finally,
@@
This implies that $[b,c,k\circ k_{1}^{\prime}]=0$, and the proof is concluded. ∎
@@
An extension $(A,f)$ in $\mathsf{SKB}$ is central with respect to the adjunction (4.1) if and only if $\operatorname{Ker}(f)\subset Z_{R}(A)$ or equivalently if $[\operatorname{Ker}(f),A]_{\mathsf{RadRng}}=0$.
@@
for the “object part” of the pullback in diagram (2.2). Let us first prove that $a+b-a-b=0$ for any $a\in\operatorname{Ker}(f)$, $b\in A$. The elements $(a,0)$ and $(b,b)$ are in $A\times_{B}A$, hence
@@
Conversely, assume now that $[\operatorname{Ker}(f),A]_{\mathsf{RadRng}}=0$. Let $(a,a+k)\in A\times_{B}A$, $(a_{1},a_{1}+k_{1})\in A\times_{B}A$ and $(a_{2},a_{2}+k_{2})\in A\times_{B}A$ be such that $k,k_{1},k_{2}\in\operatorname{Ker}(f)$. It will suffice to prove that the equality $R(s)=R(t)$ holds for all the generators of $R(A\times_{B}A)=[A\times_{B}A,A\times_{B}A]_{\mathsf{RadRng}}$. For this, observe that the assumption $\operatorname{Ker}(f)\subset Z(A_{+})$ directly implies that
@@
The arguments used in the proof of Theorem 4.13 also provide a characterization of central extensions of skew left braces with respect to the subvariety $\mathsf{BR}$ of braces [25], as we now briefly explain. The subvariety $\mathsf{BR}$ of $\mathsf{SKB}$ is determined by the additional identity $a+b-a-b=0$, since braces are precisely left skew braces of abelian type [13]. We then have the following adjunction
[…]
where $U_{B}$ is the inclusion functor and $\mathsf{br}$ its left adjoint sending a skew brace $A$ to the quotient $\mathsf{br}(A)=\frac{A}{[A,A]_{\mathsf{BR}}}$, with $[A,A]_{\mathsf{BR}}$ the ideal of $A$ generated by all the commutators $[a,b]=a+b-a-b$ (for any $a,b\in A$). One can then see that an extension $f\colon A\rightarrow B$ of skew braces is central with respect to the subvariety $\mathsf{BR}$ of braces if and only if its kernel $\operatorname{Ker}(f)$ satisfies the condition $[\operatorname{Ker}(f),A]_{\mathsf{BR}}=0$, where $[\operatorname{Ker}(f),A]_{\mathsf{BR}}$ is the ideal of $A$ generated by all the commutators of the form $[k,a]=k+a-k-a$, where $k\in\operatorname{Ker}(f),a\in A$. Indeed, by looking at the first part of the proofs of the two implications in Theorem 4.13 one realizes that the condition $[\operatorname{Ker}(f),A]_{\mathsf{BR}}=0$ is indeed necessary and sufficient for an extension to be central with respect to the adjunction (4.3). Similar results can be obtained by considering the subvarieties $\mathsf{Nil_{n}SKB}$ and $\mathsf{Sol_{n}SKB}$ of $\mathsf{SKB}$, where $\mathsf{Nil_{n}SKB}$ denotes the variety of skew left braces of $n$-nilpotent type and $\mathsf{Sol_{k}SKB}$ the variety of skew braces of $n$-solvable type ($n\geq 1$ is a positive integer). In the group case the characterizations of all these types of relative central extensions were established in [9] (Section $9$). We shall not pursue this further in the present article, but it would be interesting to describe the corresponding relative commutators of skew braces on the model of what was done in [6] in the category of groups.

### v2

@@ 1. Introduction
Skew braces form a variety of universal algebras $\mathsf{SKB}$, and generalise at the same time groups and radical rings. Concretely, given a group $(G,\cdot)$ one can give $G$ a skew brace structure by taking $\cdot$ and $\circ=\cdot$. In particular, this means that the variety $\mathsf{Grp}$ of groups is a subvariety of the variety $\mathsf{SKB}$ of skew braces. On the other hand, a radical ring $(R,+,\cdot)$ is a (not necessarily commutative) ring without unit such that the operation $a\circ b=a+a\cdot b+b$ is a group operation. In this case, the triple $(R,+,\circ)$ is a skew brace. We will recall in Section 4 a characterisation of radical rings [25] that implies that radical rings also form a subvariety $\mathsf{RadRng}$ of the variety $\mathsf{SKB}$ of skew braces. Since skew braces form a variety of $\Omega$-groups in the sense of Higgins [14], hence in particular a semi-abelian category [19] (in fact, even a strongly protomodular category [2, 3]), it is natural to look at non-abelian homology of skew braces with coefficient functors in the subvarieties $\mathsf{Grp}$ of groups, $\mathsf{Ab}$ of abelian groups, and $\mathsf{RadRng}$ of radical rings.
@@
via the classical Hopf formula, i.e. the right-hand isomorphism. The commutators $[K,F]$ and $[F,F]$ appearing in this formula are “relative” to the chosen subvariety of the variety $\mathsf{Grp}$ of groups, in this case the variety $\mathsf{Ab}$ of abelian groups, so that they could also be denoted by $K,F]_{\mathsf{Ab}}$ and $F,F]_{\mathsf{Ab}}$, respectively. The homology group $\mathsf{H}_{2}(G)$ can also be obtained by applying comonadic homology theory [1] to the variety $\mathsf{Grp}$ with respect to the abelianisation functor $\mathsf{ab}\colon\mathsf{Grp}\rightarrow\mathsf{Ab}$, so that $\mathsf{H}_{2}(G)\cong\mathsf{H}_{2}(G,\mathsf{ab})$, the right-hand side homology group now being the one arising from the comonadic approach by taking $\mathsf{ab}$ as coefficient functor:
@@
First, in Section 3, we characterize the central extensions in $\mathsf{SKB}$ corresponding to the subvariety $\mathsf{Grp}$ of groups (Proposition 3.5) that are the surjective homomorphisms $f\colon A\rightarrow B$ of skew braces such that $\operatorname{Ker}(f),A]_{\mathsf{Grp}}=0$, where $\operatorname{Ker}(f),A]_{\mathsf{Grp}}$ is the additive subgroup of $A$ generated by all the elements of the form
@@
Section 4 is devoted to the study of central extensions of skew braces relative to the subvariety $\mathsf{RadRng}$ of radical rings. For this purpose we introduce a new object, the radicalator (Definition 4.9). This is generated as a subgroup of the additive group by the elements $(a+b)\circ c-b\circ c+c-a\circ c$ and a+baba+b-a-b for all $a,b,c\in A$. The radicalator turns out to be an ideal $A,A]_{\mathsf{RadRng}}$ (Proposition 4.8) playing a similar role with respect to the reflection from $\mathsf{SKB}$ to $\mathsf{RadRng}$ to the role played by the derived subgroup in the reflection from $\mathsf{Grp}$ to $\mathsf{Ab}$. Indeed, the quotient $A\rightarrow\frac{A}{[A,A]_{\mathsf{RadRng}}}$ is the universal one transforming a skew brace $A$ into a radical ring $\frac{A}{[A,A]_{\mathsf{RadRng}}}$, and an extension $f\colon A\rightarrow B$ of skew braces is central relatively to the subvariety $\mathsf{RadRng}$ if and only if $\operatorname{Ker}(f),A]_{\mathsf{RadRng}}=0$ (Theorem 4.13). This pushes forward Rump’s connection between skew braces and radical rings [25] (see also [20]). We also observe that the central extensions of skew braces relative to the subvariety of braces (= skew braces of abelian type) can also be characterized in algebraic terms (Remark 4.14).
@@
is shown to be an ideal (Proposition 5.3), denoted by $I,A]_{\mathsf{Ab}}$ (here $a,b]_{+}$ and $a,b]_{\circ}$ are the usual group-theoretic commutators with respect to the operations ++ and $\circ$, respectively). An extension $f\colon A\rightarrow B$ of skew braces is then central with respect to $\mathsf{Ab}$ if and only if $\operatorname{Ker}(f),A]_{\mathsf{Ab}}=0$ (Proposition 5.4). Note that these central extensions of skew braces have already been studied in [23, 24, 21]. We also observe that the commutator $\operatorname{Ker}(f),A]_{\mathsf{Ab}}$ coincides with the Huq commutator defined in [15] (Proposition 5.5).
@@ 2. Preliminaries
We can introduce a binary operation $\colon A\times A\rightarrow A$ defined by $a*b=-a+a\circ b-b$ for all $a,b\in A$. By definition, a normal subgroup $I$ of $A_{+}$ is an ideal of $A$ if $a*x\in I$ and $x*a\in I$ for all $a\in A$ and $x\in I$. A normal subgroup $I$ of the group $(A,+)$ such that $I$ is a subgroup of $A_{\circ}$ and $A*I\subset I$ is called a strong left ideal.
@@
[…]
where $A\times_{B}A=\{(a_{1},a_{2})\in A\times A\,\mid\,f(a_{1})=f(a_{2})\}$, while $s$ and $t$ are the first and second projection, respectively. The definition of central extension relative to a given subvariety $\mathcal{X}$ of $\mathsf{SKB}$ (as in the work of the Fröhlich school [11, 12, 22]) is then the following:
@@ 3. The subvariety of groups
where $\mathsf{Grp}$ is the subvariety of $\mathsf{SKB}$ whose objects are skew braces with the property that the two group operations ++ and $\circ$ are equal. This subvariety is isomorphic to the category of groups, of course, and this justifies the slight abuse of notation. The functor $U_{G}$ is a full inclusion, and in order to describe its left adjoint $F$ in (3.1) the following definition will be needed:
@@
Let $A$ be a skew brace, we denote by AAA*A the subgroup of $A_{+}$ generated by the elements aba*b for all $a,b\in A$.
It is well known that AAA*A is an ideal of $A$ and it is the smallest ideal $I$ such that the quotient $A/I$ is a group.
In our situation the reflector $F:\mathsf{SKB}\to\mathsf{Grp}$ is the functor sending the skew brace $A$ to the quotient F(A)=A/(AA)F(A)=A/(A*A) by the ideal AAA*A (with obvious definition on morphisms). Note that here the ideal AAA*A is precisely the kernel $R(A)$ of the $A$-component of the unit $\eta_{A}\colon A\rightarrow A/(A*A)$ of the adjunction (3.1). This adjunction is admissible in the sense of Categorical Galois theory, since $\mathsf{Grp}$ is a subvariety of the variety $\mathsf{SKB}$ (see Section $5$ in [18]).
@@
Let $A$ be a skew brace and $I\subset A$ an ideal, we define $I,A]_{\mathsf{Grp}}$ to be the additive subgroup of $A$ generated by the set
@@
Let $A$ be a skew brace and $I\subset A$ an ideal of $A$. Then $I,A]_{\mathsf{Grp}}$ is an ideal.
@@
An extension $(A,f)$ is central with respect to the adjunction (3.1) if and only if $\operatorname{Ker}(f),A]_{\mathsf{Grp}}=0$.
@@
Since $(A,f)$ is central, by definition ab=R(s)(ab,0)=R(t)(ab,0)=0a*b=R(s)(a*b,0)=R(t)(a*b,0)=0. Similarly, one can show that ba=0b*a=0. As $b\in A$ and $a\in\operatorname{Ker}(f)$ were chosen arbitrarily, it follows that $\operatorname{Ker}(f),A]_{\mathsf{Grp}}=0$.
Assume now that $\operatorname{Ker}(f),A]_{\mathsf{Grp}}=0$. Let $(a,b)$ and $(a_{1},b_{1})$ be two elements in $A\times_{B}A$. Since $a_{1}+b_{1}\in\operatorname{Ker}(f)$ one has that $a_{1}\circ(-a_{1}+b_{1})=b_{1}$ so that $a_{1}+b_{1}=a_{1}^{\prime}\circ b_{1}$. It follows that
@@ 4. The subvariety of radical rings
$a,b,c]=(a+b)\circ c-b\circ c+c-a\circ c.$
@@
Let $A$ be a skew brace, and $I\subset A$ an ideal. We define $I,A]_{\mathsf{RadRng}}$ to be the additive subgroup of $A$ generated by the elements $[a,b,c]$, $[c,a,b]$, $[b,c,a]$ and a+baba+b-a-b for all $a\in I$ and $b,c\in A$.
Our next goal is to show that $I,A]_{\mathsf{RadRng}}$ is an ideal.
@@
If we write $J=[I,A]_{\mathsf{RadRng}}$, then $J_{+}$ is a normal subgroup of $A_{+}$. Indeed, for any $j\in J\subset I$ and $a\in A$, the element j+ajaj+a-j-a belongs to $J$, so that $J$ is stable under conjugation in $A$.
@@
To see that $\lambda_{a}(J)\subset J$, it suffices to check on the generators of $J$. It is clear that $\lambda_{d}([I_{+},A_{+}])\subset[I_{+},A_{+}]\subset J$ for all $d\in A$. Let then $d\in A$, we must show that $\overline{\lambda_{d}([a,b,c])}=0$ in the quotient group $A_{+}/J_{+}$ where $a,b,c\in A$ and at least one of these elements is in $I$.
@@
$=\overline{(d\circ a+\lambda_{d}(b)-d\circ b+d-d\circ a)\circ c-c}$
@@
$c+\lambda_{d}(b)\circ c\in I,\quad-d\circ b\circ c+d\circ c=\lambda_{d}(-b\circ c+c)\in I.$
@@
Let $A$ be a skew brace, and $I\subset A$ an ideal. The subgroup $I,A]_{\mathsf{RadRng}}$ of $A_{+}$ is an ideal of $A$.
@@
Let $J=[I,A]_{\mathsf{RadRng}}$. It is left to show that $J*A\subset J$. It suffices to show that $k\circ b-b\in J$ for all $k\in J$ and $b\in A$. Since $J\subset I$, it is sufficient to show this on the generators of $J$. Let $a\in I$ and $b,c\in A$. By Lemma 4.5,
@@
To obtain the second and third equality, we used Corollary 4.6. Assume now that $b\in I$ and $a,c,d\in A$. Then
@@
Finally, assume that $c\in I$ and $a,b,d\in A$. In this case,
@@
Therefore $k^{\prime}\circ k_{1},b,c]=0$. Similarly one sees that $b,k^{\prime}\circ k_{1},c]=0$. Finally,
@@
This implies that $b,c,k\circ k_{1}^{\prime}]=0$, and the proof is concluded. ∎
@@
An extension $(A,f)$ in $\mathsf{SKB}$ is central with respect to the adjunction (4.1) if and only if $\operatorname{Ker}(f)\subset Z_{R}(A)$ or equivalently if $\operatorname{Ker}(f),A]_{\mathsf{RadRng}}=0$.
@@
for the “object part” of the pullback in diagram (2.2). Let us first prove that a+bab=0a+b-a-b=0 for any $a\in\operatorname{Ker}(f)$, $b\in A$. The elements $(a,0)$ and $(b,b)$ are in $A\times_{B}A$, hence
@@
Conversely, assume now that $\operatorname{Ker}(f),A]_{\mathsf{RadRng}}=0$. Let $(a,a+k)\in A\times_{B}A$, $(a_{1},a_{1}+k_{1})\in A\times_{B}A$ and $(a_{2},a_{2}+k_{2})\in A\times_{B}A$ be such that $k,k_{1},k_{2}\in\operatorname{Ker}(f)$. It will suffice to prove that the equality $R(s)=R(t)$ holds for all the generators of $R(A\times_{B}A)=[A\times_{B}A,A\times_{B}A]_{\mathsf{RadRng}}$. For this, observe that the assumption $\operatorname{Ker}(f)\subset Z(A_{+})$ directly implies that
@@
The arguments used in the proof of Theorem 4.13 also provide a characterization of central extensions of skew left braces with respect to the subvariety $\mathsf{BR}$ of braces [25], as we now briefly explain. The subvariety $\mathsf{BR}$ of $\mathsf{SKB}$ is determined by the additional identity a+bab=0a+b-a-b=0, since braces are precisely left skew braces of abelian type [13]. We then have the following adjunction
(4.3) $\mathsf{BR}$$\perp$$\mathsf{SKB}$$\scriptstyle{U_{B}}$br
where $U_{B}$ is the inclusion functor and $\mathsf{br}$ its left adjoint sending a skew brace $A$ to the quotient $\mathsf{br}(A)=\frac{A}{[A,A]_{\mathsf{BR}}}$, with $A,A]_{\mathsf{BR}}$ the ideal of $A$ generated by all the commutators [a,b]=a+bab[a,b]=a+b-a-b (for any $a,b\in A$). One can then see that an extension $f\colon A\rightarrow B$ of skew braces is central with respect to the subvariety $\mathsf{BR}$ of braces if and only if its kernel $\operatorname{Ker}(f)$ satisfies the condition $\operatorname{Ker}(f),A]_{\mathsf{BR}}=0$, where $\operatorname{Ker}(f),A]_{\mathsf{BR}}$ is the ideal of $A$ generated by all the commutators of the form [k,a]=k+aka[k,a]=k+a-k-a, where $k\in\operatorname{Ker}(f),a\in A$. Indeed, by looking at the first part of the proofs of the two implications in Theorem 4.13 one realizes that the condition $\operatorname{Ker}(f),A]_{\mathsf{BR}}=0$ is indeed necessary and sufficient for an extension to be central with respect to the adjunction (4.3). Similar results can be obtained by considering the subvarieties $\mathsf{Nil_{n}SKB}$ and $\mathsf{Sol_{n}SKB}$ of $\mathsf{SKB}$, where $\mathsf{Nil_{n}SKB}$ denotes the variety of skew left braces of $n$-nilpotent type and $\mathsf{Sol_{k}SKB}$ the variety of skew braces of $n$-solvable type ($n\geq 1$ is a positive integer). In the group case the characterizations of all these types of relative central extensions were established in [9] (Section $9$). We shall not pursue this further in the present article, but it would be interesting to describe the corresponding relative commutators of skew braces on the model of what was done in [6] in the category of groups.